# Reinforced MLLM: A Survey on RL-Based Reasoning in Multimodal Large Language Models

### v1

@@ 8. Conclusion
In this survey, we presented a comprehensive overview of reinforcement learning (RL)-based reasoning methods in Multimodal Large Language Models (MLLMs). We analyzed recent progress in RL optimization strategies, categorized value-free and value-based approaches, and discussed reward mechanism designs that enable MLLMs to reason across text, vision, audio, and video modalities. By summarizing benchmark datasets, evaluation protocols, and practical applications, we provided a structured understanding of how RL is shaping the development of multimodal reasoning. While significant advancements have been made, RL-based multimodal reasoning still faces critical challenges. Current methods often rely on text-centric reasoning paths, underutilize multimodal information, and depend heavily on simple, verifiable reward designs. Furthermore, issues such as sparse and delayed rewards, inefficient exploration strategies, and limited generalization to open-ended tasks remain key obstacles. Looking ahead, future research should explore hierarchical and structured reward models, dynamic cross-modal Chain-of-Thought generation, and lightweight, scalable RL frameworks. Addressing these challenges is crucial for enabling MLLMs to perform more robust, interpretable, and generalized reasoning across real-world multimodal tasks. We hope this survey will serve as a valuable reference for researchers and practitioners, and inspire further innovation in this rapidly evolving field.

### v2

@@ 8. Conclusion
In this survey, we presented a comprehensive overview of RL-based reasoning methods in MLLMs. We analyzed recent progress in RL optimization strategies, categorized value-model-free and value-model-based approaches, and discussed reward mechanism designs that enable MLLMs to reason across text, vision, audio, and video modalities. By summarizing benchmark datasets, evaluation protocols, and practical applications, we provided a structured understanding of how RL is shaping the development of multimodal reasoning. While significant advancements have been made, RL-based multimodal reasoning still faces critical challenges. Current methods often rely on text-centric reasoning paths, underutilize multimodal information, and depend heavily on simple, verifiable reward designs. Furthermore, issues such as sparse and delayed rewards, inefficient exploration strategies, and limited generalization to open-ended tasks remain key obstacles. Looking ahead, future research should explore hierarchical and structured reward models, dynamic cross-modal CoT generation, and lightweight, scalable RL frameworks. Addressing these challenges is crucial for enabling MLLMs to perform more robust, interpretable, and generalized reasoning across real-world multimodal tasks. We hope this survey will serve as a valuable reference for researchers and practitioners, and inspire further innovation in this rapidly evolving field.